# ZoFia: Zero-Shot Fake News Detection with Entity-Guided Retrieval and Multi-LLM Interaction

### v1

@@ Abstract
The rapid spread of fake news threatens social stability and public trust, rendering its detection an imperative research priority. Although large language models (LLMs) excel at numerous natural language processing tasks with their remarkable contextual understanding and extensive prior knowledge, the time-bounded knowledge coverage and tendency for generating hallucination content reduce their reliability when handling fast-evolving news streams. Furthermore, models trained on existing static datasets also often lack the generalization needed for emerging news topics. To address these challenges, we propose ZoFia, a novel two-stage zero-shot fake news detection framework. First, we introduce Hierarchical Salience to quantify the importance of entities in the news content, and propose the SC-MMR algorithm to effectively select an informative and diverse set of keywords that serve as queries for retrieving up-to-date external evidence. Subsequently, a multi LLM interactive system, in which each agent assumes a distinct role, performs multi-view collaborative analysis and adversarial debate over the news text and its related information, and finally produces an interpretable and robust judgment. Comprehensive experiments on two public datasets demonstrate that ZoFia obviously outperforms existing zero-shot baselines and most of few-shot methods. Our codes will be open-sourced to facilitate related communities.
@@
Lvhua Wu1,2thanks: Equal contribution., Xuefeng Jiang1,211footnotemark: 1, Sheng Sun1, Tian Wen1, Yuwei Wang1, Min Liu1,2thanks: Corresponding author. 1Institute of Computing Technology, Chinese Academy of Sciences 2University of Chinese Academy of Sciences {wulvhua24s,jiangxuefeng21b,liumin}@ict.ac.cn
@@ 1 Introduction
The rapid spread of fake news through social networks has become prevalent, posing severe threats to key domains such as politics Fisher et al. (2016), economy Bakir and McStay (2018), and livelihood Zhou and Zafarani (2020). The swift advancement of generative models further exacerbates this concern Chen and Shu (2023); Nan et al. (2024); Li et al. (2025, 2024). In this context, developing effective, efficient and interpretable fake news detection methods is imperative.
Fake news detection is formally defined as a binary classification problem. Prior to the rise of Large Language Models (LLMs), related research is predominantly led by training deep models via supervised learning Monti et al. (2019); Kaliyar et al. (2021). LLMs exhibit significant potential in information forensics due to their broad knowledge and strong contextual understanding acquired through pretraining on massive corpora Su et al. (2023). Current research approaches utilizing LLMs for detection generally fall into two categories. The first category employs LLMs as zero-shot reasoners, directly guiding models to judge and explain news authenticity through meticulously designed prompt engineering Jiang et al. (2022) or instruction tuning Pavlyshenko (2023). The second category treats LLMs as auxiliary tools, leveraging their powerful generative capabilities for information augmentation to enhance the performance and robustness of downstream compact classifiers Hu et al. (2024a).
However, directly applying LLMs to fake news detection poses significant challenges. The main problem is that LLMs have a knowledge cutoff Cheng et al. (2024), which means that an LLM’s internal knowledge becomes static after training, limiting its capacity to adapt to dynamic news streams. Furthermore, LLMs are prone to hallucinations Ji et al. (2023), frequently producing plausible but factually incorrect content, which directly undermines its reliability as a misinformation detector. More seriously, existing supervised methods generalize poorly to emerging topics Hoy and Koulouri (2022), and a single LLM used for zero-shot judgment often produces shallow reasoning, leading to superficial analysis.
To address these challenges systematically, we propose ZoFia, a training-free zero-shot framework for fake news detection. ZoFia decouples the complex detection task into two stages. First, it performs entity guided retrieval to integrate the LLM’s internal knowledge with real time external information and builds a comprehensive Multi-Source Information Matrix. Next, an multi-LLM interaction system conducts multidimensional analysis and adversarial debate over this matrix to deliver an explainable and robust judgment. The framework aims to compensate for the knowledge limits of LLMs while fully leveraging their inherent abilities of reasoning and interaction.
Our main contributions can be summarized as follows:
@@
  • We propose ZoFia, a zero-shot fake news detection framework that combines entity-guided retrieval with multi-agent collaboration, effectively addressing LLMs’ knowledge cutoff and hallucinations.
  • We propose two novel algorithms, Hierarchical Salience and SC-MMR, to achieve accurate keyword extraction for news.
@@
  • Experiments conducted on two public real-world datasets demonstrate that our proposed framework outperforms existing zero-shot models and most of few-shot models.
  • Code will be open-sourced to ensure the reproduction and facilitate related communities.
@@ 2 Related Work
2.1 LLM-based Fake News Detection
Using LLMs for fake news detection is a recent research trend, and current studies explores various capabilities of LLM. Hu et al. Hu et al. (2024a) position LLMs as auxiliary tools and propose that while LLMs cannot directly serve as detectors, they can provide multi-perspective guidance signals to smaller models. GenFEND Nan et al. (2024) focuses on information augmentation, which utilizes LLMs to generate diverse comments simulating different user groups to enrich related information. To address LLM knowledge limitations, Liu et al. Liu et al. (2024) combine internal sample retrieval with external real-time information retrieval, alleviating issues of vague and scarce information.
However, these methods seldom deeply explore other capabilities of LLMs, including inherent reasoning and role-playing. Distinct from prior studies, our framework guides LLMs to play different roles, collaboratively analyzing news from multiple expert perspectives and achieving better and more robust zero-shot detection.
2.2 Multi-agent Collaboration
Multi-agent collaboration as a new paradigm for enhancing LLM capabilities remains in its early stages. Its core concept involves integrating multiple agents to strengthen processing capacity for complex tasks. As pioneering work, Chateval Chan et al. (2023) demonstrated through agent team-based evaluation that this approach simultaneously improves robustness and accuracy in natural language generation. COLA Lan et al. (2024) applied this concept to zero-shot stance detection, assigning different roles to LLMs and combining a debate framework to enhance reasoning, yet failed to effectively mitigate hallucination problems. TruEDebate Liu et al. (2025) introduced multi-stage structured debate into fake news detection, but its implementation revealed a tendency toward premature convergence of viewpoints.
Our work uses a role-playing multi-agent paradigm, but also incorporates external knowledge retrieval and independent modular analysis to reduce hallucination and ensure diversity and independence of debate arguments.
3 Entity-Guided Retrieval
The efficiency of LLMs is constrained by the knowledge cutoff from training data Cheng et al. (2024). This limitation becomes especially apparent when processing rapidly changing real-time news streams, as misinformation usually propagates at rates and patterns that outpace the models’ training cycles Yi et al. (2025). Therefore, enhancing LLMs’ knowledge through external information retrieval is necessary.
We design the entity-guided retrieval stage to acquire real-time and reliable external knowledge for subsequent use by LLM experts. The framework consists of 4 sequential modules. The first 3 modules precisely extract and filter out a set of critical entities from the original news text as the keywords. These keywords are then used as queries for the final retrieval module, which retrieves from both Open Web and Wikipedia. This stage provides LLMs with the essential and up-to-date factual grounding and contextual information.
@@ 3.1 Entity Extractor
This module first uses a pre-trained BERT-NER Tjong Kim Sang and De Meulder (2003) model to perform named entity recognition (NER) on the news text. This process can be expressed as:
@@ 3.2 Salience Scorer
This module aims to accurately quantify the importance of key entities in news texts, i.e. Entity Salience (ES) Dunietz and Gillick (2014). Although the prior study Bullough et al. (2024) has demonstrated that bi-encoder architectures can efficiently estimate entity salience, their performance is often limited by the semantic dilution Hou et al. (2021) problem. In lengthy or unevenly distributed texts, the core semantics of key entities are diluted by minor content, resulting in their true importance being severely underestimated.
@@ 3.3 Keyword Selector
$\mathcal{M}_{\text{SBERT}}$: SBERT encoder for entity embeddings;
$\gamma$: Decay factor for termination criterion
Output: $\mathcal{U}_{\text{selected}}$: Final set of selected entities
if $\mathcal{U}_{\text{candidate}}=\emptyset$ then
 return $\mathcal{U}_{\text{selected}}$
$\text{MMR}_{\text{prev}}^{*}\leftarrow 1.0$
while $\mathcal{U}_{\text{candidate}}\neq\emptyset$ do
 $k\leftarrow|\mathcal{U}_{\text{selected}}|$
 $\lambda_{k}\leftarrow\max(1.0\,,\,1.2-e^{0.2k-1.5}/e)$
 foreach $U_{i}\in\mathcal{U}_{\text{candidate}}$ do
    if $\text{MMR}_{\text{curr}}>\text{MMR}_{\text{best}}$ then
 if $\text{MMR}_{\text{best}}\leq\gamma\cdot\text{MMR}_{\text{prev}}^{*}$ then
    break
 $\text{MMR}_{\text{prev}}^{*}\leftarrow\text{MMR}_{\text{best}}$
return $\mathcal{U}_{\text{selected}}$
This module aims to select an informative subset of keywords from the candidate entity set $\mathcal{U}_{\text{selected}}$ to mitigate the query drift Carpineto and Romano (2012) problem. However, this process faces two practical challenges. First, The hierarchical salience score $\mathcal{S}_{\text{hier}}(U_{i})$ is highly sensitive to context granularity, so a fixed screening threshold is not effective. Second, coreference in news introduces high–scoring entities with repeated semantics, which harms the diversity of the keyword set.
To optimize relevance and diversity under these constraints, We propose an improved MMR Carbonell and Goldstein (1998) algorithm, namely Salience-Calibrated MMR (SC-MMR) as shown in Algorithm 1. At the $k$-th iteration, SC-MMR evaluates the score of a candidate entity $U_{i}$ by
@@
The core innovation of SC-MMR is a dynamic decaying weight $\lambda_{k}$. At the early stage, the weight places more emphasis on an entity’s own salience. As the number of selected entities $k$ increases, the weight adaptively decreases for high diversity and low redundancy. The computation of $\lambda_{k}$ is
$\lambda_{k}=\max\left[0.1\,,\,1.2-\frac{\exp(0.2k-0.5)}{e}\right].$ (5)
We further replace termination strategies that rely on a fixed score with a dynamic criterion based on the relative change of the MMR score. The iteration continues only when the next best candidate $\mathcal{U}_{k+1}^{*}$ satisfies $\text{MMR}(\mathcal{U}_{k+1}^{*})>\gamma\cdot\text{MMR}(\mathcal{U}_{k}^{*})$, where $\gamma$ is a decay factor. It lets the algorithm determine the stopping point according to score distribution and prevents a decline in the quality of entities caused by diminishing marginal utility.
@@ 3.4 Information Retrieval
This module uses the keyword set $\mathcal{K}=\{k_{1},k_{2},\ldots,k_{N}\}$ distilled in previous modules to build a comprehensive external knowledge base $\mathcal{E}$, serving as supplementary context for the subsequent detection stage. We implement a dual source retrieval that gathers information from the open web and Wikipedia in parallel, integrating the timeliness of the former with the authoritativeness of the latter.
Open Web. We compose all keywords into an aggregated query for the open web $Q_{\text{web}}$ with the following logic:
$Q_{\text{web}}=\left(\bigwedge_{i=1}^{N}k_{i}\right)\land\neg\texttt{news}\land\neg\texttt{wikipedia}.$ (6)
The operator $\land$ requires that results relate to all keywords, which enables strict cross-keyword verification. We explicitly exclude news and wikipedia sources. The former avoids retrieving the duplicate reports to prevent source contamination Deng et al. (2023), and the latter prevents redundancy with the subsequent Wikipedia retrieval. For each returned page, we only extract its summary and search snippet as the raw corpus.
Wikipedia. The summary section of a Wikipedia entry usually provides the most precise definition for an entity. However, a single keyword $k_{i}$ often corresponds to multiple Wikipedia entries, which introduces semantic ambiguity. We build a context-aware disambiguation mechanism that uses the original local context $\mathcal{C}(U_{i})$ of keyword $k_{i}$ in the news text to perform accurate matching.
@@
4 Multi-LLM Interaction
Entity-guided retrieval supplies the framework with timely external information, which together with the prior knowledge of LLMs builds a Multi-Source Information Matrix. This stage employs a dual-state multi-LLM interaction system that fully exploits this matrix to mitigate hallucinations Feng et al. (2024) and ensure accurate detection.
This stage operates in two orthogonal interaction states. LLM Collaboration performs parallel analyses across multiple agents to reduce inferential variance. LLM Debate and Judgment introduces a adversarial debate to reduce systemic bias. These two states form a complete reasoning chain from mining divergent evidences to making a convergent judgment.
@@ 4.1 LLM Collaboration
LLM Collaboration state aims to reduce inferential variance. Through parallel analysis by multiple agents, it transforms Multi-Source Information Matrix from the previous stage into a structured evidence pool, which provides a stable decision basis for the subsequent adversarial debate. As illustrated in Figure 2, the matrix systematically integrates 4 information sources that are orthogonal and complementary in evidence distribution:
  • In-news Information: It comes from the original text of the target news and provides the basic core content.
  • Out-of-news Information: It comes from Open Web retrieval and offers the broadest instant facts.
  • LLM internal Knowledge: It comes from the model’s prior knowledge and provides generalized common sense.
  • LLM external Knowledge: It comes from Wikipedia retrieval and provides the most precise summary for the core entities.
@@ 4.1.1 Linguist
  • Emotion: Proportion of affective terms in the text and the headline, and the degree of incendiary tone.
@@
To maintain objective independence, each dimension is evaluated in an isolated session. For each dimension, LLM explicitly indicates whether it reflects the news is real or fake. This agent uses 2 quadrants of Multi-source Information Matrix: LLM internal knowledge and in-news information from the original news text.

### v3

@@ Abstract
The rapid spread of fake news threatens social stability and public trust, highlighting the urgent need for its effective detection. Although large language models (LLMs) show potential in fake news detection, they are limited by knowledge cutoff and easily generate factual hallucinations when handling time-sensitive news. Furthermore, the thinking of a single LLM easily falls into early stance locking and confirmation bias, making it hard to handle both content reasoning and fact checking simultaneously. To address these challenges, we propose ZoFia, a two-stage zero-shot fake news detection framework. In the first retrieval stage, we propose novel Hierarchical Salience and Salience-Calibrated Minimum Marginal Relevance (SC-MMR) algorithm to extract core entities accurately, which drive dual-source retrieval to overcome knowledge and evidence gaps. In the subsequent stage, a multi-agent system conducts multi-perspective reasoning and verification in parallel and achieves an explainable and robust result via adversarial debate. Comprehensive experiments on two public datasets show that ZoFia outperforms existing zero-shot baselines and even most few-shot methods. Our code has been open-sourced to facilitate the research community at https://github.com/SakiRinn/ZoFia.
@@
Lvhua Wu1,2thanks: Equal contribution., Xuefeng Jiang1,211footnotemark: 1, Sheng Sun1, Yan Lei1,2, Tian Wen1,2, Yuwei Wang1, Min Liu1,2thanks: Corresponding author. 1Institute of Computing Technology, Chinese Academy of Sciences 2University of Chinese Academy of Sciences {wulvhua24s,jiangxuefeng21b,liumin}@ict.ac.cn
@@ 1 Introduction
The rapid spread of fake news through social networks has become prevalent, posing severe threats to key domains such as politics Fisher et al. (2016), economy Bakir and McStay (2018), and livelihood Zhou and Zafarani (2020). The swift advancement of generative models further exacerbates this concern Chen and Shu (2023); Nan et al. (2024); Li et al. (2025, 2024a). In this context, developing effective, efficient and interpretable fake news detection methods is indispensable.
Early studies mainly rely on supervised learning Monti et al. (2019); Kaliyar et al. (2021) to train detection models, but their dependence on large-scale labeled data makes it hard to adapt to emerging news topics Hoy and Koulouri (2022). Large language models (LLMs), with broad pretraining knowledge and strong contextual understanding Su et al. (2023), significantly advance few-shot and zero-shot methods, providing new opportunities to overcome this bottleneck. In few-shot methods, LLMs either serve as auxiliary tools to perform data augmentation for downstream classifiers Hu et al. (2024a), or act directly as detectors through prompt learning Jiang et al. (2022) and instruction tuning Pavlyshenko (2023), but they still do not fully escape reliance on training data. By contrast, zero-shot methods that do not require labeled samples are a more promising paradigm, which mainly guide LLM reasoning by context engineering Zhang and Gao (2023) and agentic architectures Li et al. (2024b).
However, zero-shot methods face reliability challenges. On the one hand, the internal knowledge of LLMs is limited by knowledge cutoff Cheng et al. (2024). Without the external information, they easily produce factual hallucinations Ji et al. (2023) when handling dynamic news events. On the other hand, LLMs tend to lock their stance early Echterhoff et al. (2024), where confirmation bias Wan et al. (2025) drives subsequent reasoning to merely rationalize the initial judgment Wang et al. (2025). Consequently, when external retrieval is introduced into a single LLM, it easily takes a cognitive shortcut that substitutes whether information can be retrieved for factual veracity, which seriously weakens content reasoning on the original news. This effect becomes much stronger when the retrieved evidence is irrelevant or contradictory Tan et al. (2024). Therefore, we observe that the inherent reasoning defects of a single LLM prevent it from performing both news content reasoning and external evidence verification well simultaneously.
We argue that content reasoning and fact checking should be decoupled and reliable judgment can be achieved through multi-agent interaction. Based on this motivation, we propose ZoFia, a two-stage zero-shot framework for fake news detection. In the first entity-guided retrieval stage, we introduce Hierarchical Salience to address semantic dilution Hou et al. (2021) of news entities, which fully leverages global and local semantics to score entity salience, and design a novel Salience-Calibrated Minimum Marginal Relevance (SC-MMR) algorithm to accurately extract core entities. These entities then drive dual-source retrieval from Wikipedia and Open Web to mitigate knowledge cutoff and effectively suppress hallucinations. In the subsequent multi-LLM interaction stage, ZoFia assigns agents to perform content reasoning and fact checking in parallel and breaks stance locking of a single LLM via adversarial debate, ensuring the robustness and interpretability of the final verdict.
To sum up, our main contributions are outlined as follows:
@@
  • We propose ZoFia, a retrieval-augmented multi-agent zero-shot framework for fake news detection that effectively overcomes the inherent reasoning flaws of a single LLM.
  • We introduce a novel granularity-aware metric Hierarchical Salience and design SC-MMR algorithm to extract news entities accurately for efficient retrieval.
@@
  • Comprehensive experiments on two public datasets demonstrate that ZoFia outperforms existing zero-shot baselines and even most few-shot methods.
  • We open-source our code to facilitate the research community at https://github.com/SakiRinn/ZoFia.
@@ 2 Related Work
LLM in Fake News Detection. The application of Large Language Models (LLMs) has become a research frontier in fake news detection, primarily divided into few-shot and zero-shot paradigms. In few-shot methods, researchers often use LLMs as auxiliary tools for data augmentation Hu et al. (2024a); Nan et al. (2024), or as detectors via post-training Jiang et al. (2022); Pavlyshenko (2023). However, these methods fail to completely eliminate dependence on labeled data. Zero-shot methods directly guide model reasoning through context engineering Zhang and Gao (2023) and agentic architectures Li et al. (2024b); Liu et al. (2024), making judgments without labeled samples. Nevertheless, these methods fail to address the confirmation bias arising from a single reasoning chain. We decouple retrieval from reasoning via a multi-role multi-agent system and finally aggregate all evidence for judgment, achieving more comprehensive and robust zero-shot discrimination.
Multi-Agent System. Multi-Agent Systems (MAS) have emerged as an effective paradigm to enhance LLMs for complex tasks. The pioneering work Chateval Chan et al. (2023) demonstrates that MAS improves both the robustness and accuracy of generation tasks. Subsequent studies introduce this paradigm to reasoning tasks. For instance, COLA Lan et al. (2024) designs a collaboration framework for stance detection, but it remains limited to analyzing the original text. TruEDebate Liu et al. (2025) applies structured debate to fake news detection, but its implementation tends to cause premature consensus convergence. Recent agentic systems Zhang et al. (2026) also integrate reasoning more tightly with external knowledge access. Our ZoFia introduces external information retrieval and independent modular analysis. This design aims to fundamentally mitigate knowledge cutoff and ensure the diversity and independence of arguments and analyses.
3 Stage 1: Entity-Guided Retrieval
This stage aims to acquire reliable external knowledge and instant factual evidence for the subsequent multi-agent system. It consists of four sequential modules. The first three modules precisely extract a set of core entities from the original news. These entities serve as keywords and are concatenated into a query for the final retrieval module, which retrieves from Open Web and Wikipedia.
@@ 3.1 Entity Extractor
This module first uses a pre-trained BERT-NER Tjong Kim Sang and De Meulder (2003) model to perform named entity recognition (NER) on the news text. We adopt a lightweight BERT-NER model to obtain stable span-level confidence scores for dynamic filtering. This process can be expressed as:
@@ 3.2 Salience Scorer
This module aims to accurately quantify the importance of news entities, namely Entity Salience (ES) Dunietz and Gillick (2014). Although the prior study Bullough et al. (2024) has demonstrated that bi-encoder architectures can efficiently estimate entity salience, their performance is often limited by semantic dilution Hou et al. (2021) problem, which leads to severe underestimation of the importance of key entities.
@@ 3.3 Keyword Selector
This module aims to select an informative subset of keywords from the candidate entity set $\mathcal{U}_{\text{selected}}$ to mitigate the query drift Carpineto and Romano (2012) problem. However, this process faces two practical challenges. First, the hierarchical salience score $\mathcal{S}_{\text{hier}}(U_{i})$ is highly sensitive to context granularity, so a fixed screening threshold is not effective. Second, coreference in news introduces high-scoring entities with repeated semantics, which harms the diversity of the keyword set.
To optimize relevance and diversity under these constraints, we propose an improved MMR Carbonell and Goldstein (1998) algorithm, namely Salience-Calibrated MMR (SC-MMR). At the $k$-th iteration, SC-MMR evaluates the score of a candidate entity $U_{i}$ by
@@
The key innovation of SC-MMR is to introduce a weight schedule $\lambda_{k}$ that changes with the number of selected keywords k, so that the focus gradually shifts from relevance to diversity. We adopt an annealing schedule with a lower bound:
$\lambda_{k}=\max(\lambda_{\min},\lambda_{\max}-\exp(\alpha\cdot k-\beta)).$ (5)
This form ensures that $\lambda_{k}$ decreases monotonically with $k$ while retaining a non-zero salience weight via $\lambda_{\min}$ to prevent the diversity term from fully dominating.
We further introduce a dynamic termination rule based on the relative change of the MMR score. The iteration continues only when the next best candidate $\mathcal{U}_{k+1}^{*}$ satisfies $\text{MMR}(\mathcal{U}_{k+1}^{*})>\gamma\cdot\text{MMR}(\mathcal{U}_{k}^{*})$, where $\gamma$ is a decay factor. This design prevents the decline in entity quality caused by diminishing marginal utility.
@@ 3.4 Information Retrieval
This module uses the keyword set $\mathcal{K}=\{k_{1},k_{2},\ldots,k_{N}\}$ distilled in previous modules to build a comprehensive external knowledge base $\mathcal{E}$, serving as supplementary context for the subsequent detection stage. We implement a dual-source retrieval that gathers information from the Open Web and Wikipedia in parallel, ensuring both timeliness and authority.
Retrieval from Open Web. We compose all keywords into an aggregated query for the Open Web $Q_{\text{web}}$ with the following logic:
$Q_{\text{web}}=\left(\bigwedge_{i=1}^{N}k_{i}\right)$ (6)
The operator $\land$ requires that results relate to all keywords, which enables strict cross-keyword verification. We explicitly exclude news and Wikipedia sources. The former avoids retrieving duplicate reports to prevent source contamination Deng et al. (2023), and the latter prevents redundancy with the subsequent Wikipedia retrieval. For each returned page, we only extract its summary and search snippet as the raw corpus.
Retrieval from Wikipedia. The summary section of a Wikipedia entry usually provides the most precise definition for an entity. However, a single keyword $k_{i}$ often corresponds to multiple Wikipedia entries, which introduces semantic ambiguity. We build a context-aware disambiguation mechanism that uses the original local context $\mathcal{C}(U_{i})$ of keyword $k_{i}$ in the news text to perform accurate matching.
@@
4 Stage 2: Multi-LLM Interaction
The external information provided by entity-guided retrieval and the prior knowledge of LLMs builds a Multi-Source Information Matrix. This stage employs a dual-state multi-LLM system that fully exploits this matrix to perform parallel and multi-perspective content reasoning and claim verification, which are finally aggregated to reach a robust and interpretable judgment.
This stage operates in two orthogonal interaction states. LLM Collaboration performs parallel analyses across multiple agents to reduce inferential variance. LLM Debate and Judgment introduces an adversarial debate to reduce systemic bias. These two states form a complete reasoning chain from mining divergent evidence to making a convergent judgment.
@@ 4.1 LLM Collaboration
LLM Collaboration state aims to reduce inferential variance. Through parallel analysis by multiple agents, it transforms the Multi-Source Information Matrix from the previous stage into a structured evidence pool, which provides a stable decision basis for the subsequent adversarial debate. As illustrated in Figure 2, the matrix systematically integrates 4 information sources that are orthogonal and complementary in evidence distribution:
  • In-news Information: Comes from the original text of the target news and provides the basic core content.
  • Out-of-news Information: Comes from Open Web retrieval and provides the most timely and broadest materials.
  • LLM internal Knowledge: Comes from the model’s prior knowledge and provides generalized common sense.
  • LLM external Knowledge: Comes from Wikipedia retrieval and provides the most precise summary for the core entities.
@@ 4.1.1 Linguist
  • Emotion: Affective terms in the text and the headline, and the degree of incendiary tone.
@@
To maintain objective independence, each dimension is evaluated in an isolated session. For each dimension, the LLM explicitly indicates whether it reflects the news is real or fake. It uses 2 quadrants of the Multi-Source Information Matrix: LLM internal knowledge and in-news information.